# Drivetrain simulation using variational autoencoders

### v1

@@ Abstract
This work proposes variational autoencoders (VAEs) to predict a vehicle’s jerk from a given torque demand, addressing the limitations of sparse real-world datasets. Specifically, we implement unconditional and conditional VAEs to generate jerk signals that integrate features from different drivetrain scenarios. The VAEs are trained on experimental data collected from two variants of a fully electric SUV, which differ in maximum torque delivery and drivetrain configuration. New meaningful jerk signals are generated within an engineering context through the interpretation of the VAE’s latent space. A performance comparison with baseline physics-based and hybrid models confirms the effectiveness of the VAEs. We show that VAEs bypass the need for exhaustive manual system parametrization while maintaining physical plausibility by conditioning data generation on specific inputs.
@@
Impact Statement
This study demonstrates how generative AI, specifically VAEs, can address incomplete datasets in drivetrain simulation. By augmenting real-world measurements, unconditional VAEs generate physically meaningful jerk signals without prior system knowledge, while conditional VAEs enable the generation of torque-specific signals. This reduces the reliance on costly real-world testing and extensive manual system parameterization, accelerating the design and validation of drivetrain models. The findings presented in this work can position generative AI not only as a powerful tool for augmenting drivetrain datasets, but also as a complementary methodology for simulating complex driving scenarios, thereby driving innovation in vehicle performance by reducing the need for on-road vehicle testing.
@@ 1 Introduction
In automotive engineering, engineers must rely on datasets to accurately parameterize and validate the generalization capabilities of drivetrain simulation models. Ideally, these datasets should cover the variance of drivetrain hardware characteristics across the vehicle fleet and the broad variety of driving situations. However, the issue is that measurement collection in the real world can be very expensive [1]. Consequently, simulation engineers often have to work with sparse datasets that, for example, do not fully capture the variance in the jerk signal behavior across the vehicle fleet. Vehicles with different mileage or drivetrain components with slightly different clearances can exhibit significant differences in their jerk behavior. A sparse or poorly distributed dataset, therefore, limits the generalization ability of a drivetrain model.
@@
To mitigate this issue, synthetic measurements can be generated using non-generative methods that rely on models and their corresponding inputs to simulate new signals of interest [2]. Such a supervised approach to data augmentation can be greatly limited if the required input data for a physics-based or data-driven model are missing in the recorded measurement files or cannot be measured at all [3]. Another major drawback of this conventional approach is that it requires drivetrain engineers to deeply understand the underlying physical phenomena to determine which input data are necessary for a non-generative model to simulate additional valuable signals [4]. Furthermore, the more physically relevant input information fed into the model for precise simulation results, the more accurate the simulation will likely be. This input-output imbalance necessitates a large number of input parameters (like torque and speed of the electric machine) to simulate a single particular output, such as vehicle acceleration.
@@
These limitations can be tackled with generative artificial intelligence (AI) [5, 6]. One option is to adopt unconditional architectures such as Variational Autoencoders (VAEs) [7, 8], which can be trained to generate new signals solely from a limited amount of available signals collected from hardware models, simulations, or a real system and does not require additional parametrization with respect to the system configuration. This self-supervised approach to data augmentation is purely data-driven. It eliminates the need for detailed domain expertise regarding the underlying physical principles of drivetrain simulation. While unconditional data augmentation can be highly advantageous for drivetrain engineers, it is also important to consider the value of conditional AI, for example, conditional VAEs (CVAEs) [9, 10, 11]. In this case, the generation of synthetic data is additionally conditioned on other inputs, similar to supervised data augmentation. This can significantly minimize the need for real-world on-road testing, making generative AI a powerful tool for addressing incomplete datasets.
@@
The outline of this paper is as follows. First, we introduce the drivetrain simulation problem in Section 2 and explain the different possible approaches to address it. Second, we present the theoretical background for variational autoencoders in Section 3 and extensions that yield conditional VAEs. Section 4 presents the dataset and the challenges it imposes for VAE training. We then train unconditional and conditional VAEs to address the problem of jerk signal generation and measure their performance using different metrics. We also compare the performance of the trained VAEs with respect to a physics-based and a hybrid model. The manuscript concludes in Section 5.
@@ 2 Drivetrain simulation problem
A key performance indicator for quantifying vehicle drivability is the jerk, defined as the rate of change of acceleration. During abrupt positive changes in acceleration, known as tip-in load changes, the torsion in the driveshaft fluctuates, stimulating vibrations in the natural frequency range of the drivetrain (1 to 20 Hz) [15]. These longitudinal oscillations negatively affect vehicle dynamics, leading to reduced drivability performance and driver discomfort [15, 16]. Although this issue is present in vehicles with combustion engines, it becomes more pronounced in electric vehicles due to their higher torque gradients [17]. The primary control variable for managing the jerk is torque demand, although it is also influenced by drivetrain parameters such as gear ratio, vehicle mass, and tire dynamics. Furthermore, parameters such as electric machine (ELM) speed and road conditions–referred to as driving scenario parameters–significantly impact jerk behavior.
@@
However, the accuracy of these models comes at the cost of complexity. Detailed hardware models, which have been iteratively refined by experts over time, require deep domain knowledge and significant experience. Although effective, these traditional models are not always suitable for the fast-paced and disruptive development cycles of modern electric vehicles, where rapid iteration and efficiency are critical [18].
@@
Hybrid models combine the strengths of both data-driven and physics-based approaches [22, 36, 37, 38]. In the context of drivetrain simulation, they typically begin with a simplified hardware model that captures the most critical physical phenomena. This ensures that the essential dynamics are grounded in physics. A data-driven correction term is added to account for any modeling discrepancies or to refine the model [39].
@@ 3.1 Transformation between time domain jerk signal and spectrogram
where $\operatorname{Re}(S(f,t))$ and $\operatorname{Im}(S(f,t))$ are the real and imaginary parts of $S(f,t)$, respectively. To compress the dynamic range of the magnitude spectrogram, a logarithmic scaling is applied,
@@ 3.2 Variational Autoencoder (VAE)
The hidden random generative process of $\mathbf{x}$ consists of two steps: First, the latent variable $\mathbf{z}$ is sampled from some prior distribution $p(\mathbf{z})$, i.e. $\mathbf{z}\sim p(\mathbf{z})$. Second, a value $\mathbf{x}$ is generated by some conditional distribution $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$, i.e. $\mathbf{x}\sim p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$. We aim to find $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ (the parameters $\bm{\theta}$), such that we can sample from $p(\mathbf{z})$ and with high probability, $\mathbf{x}\sim p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ follows the same distribution as the $\mathbf{x}$ in our dataset [9]. From a mathematical point of view, this can be achieved by maximizing the marginal log-likelihood of $\mathbf{x}$,
$\log p_{\bm{\theta}}(\mathbf{x})=\log\int p_{\bm{\theta}}(\mathbf{x},\mathbf{z})\,\text{d}\mathbf{z}=\log\int p(\mathbf{z})\,p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})\,\text{d}\mathbf{z}.$ (3.5)
This marginal likelihood $p_{\bm{\theta}}(\mathbf{x})$ is challenging to compute because:
  • The latent variable $\mathbf{z}$ may be high-dimensional (Monte Carlo-based integration becomes almost impossible).
  • The likelihood $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ is commonly quite complicated, as it is parameterized through a complex model (e.g., a neural network), making the integral intractable.
@@
$(\bm{\mu}_{\bm{\phi}}(\mathbf{x}),\bm{\sigma}_{\bm{\phi}}(\mathbf{x}))=\mathcal{NN}_{\text{enc}}(\mathbf{x};\bm{\phi}),$ (3.8)
where $\mathcal{NN}_{\text{enc}}$ is the encoder neural network parametrized in $\bm{\phi}$, that branches into the two outputs $(\bm{\mu}_{\bm{\phi}}(\mathbf{x}),\bm{\sigma}_{\bm{\phi}}(\mathbf{x}))$. The parameter set $\bm{\phi}$ includes all weights and biases necessary to produce both $\bm{\mu}_{\bm{\phi}}(\mathbf{x})$ and $\bm{\sigma}_{\bm{\phi}}(\mathbf{x})$. The encoder bridges the data space and the latent space.
Decoder: It approximates the likelihood $p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z})$ with a neural network $\mathcal{NN}_{\mathrm{dec}}(\mathbf{z};\bm{\theta})$, parameterized by $\bm{\theta}$, to output the mean of a Gaussian distribution
$p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z})=\mathcal{N}\bigl{(}{\mathbf{x}}\mid\mathcal{NN}_{\mathrm{dec}}(\mathbf{z};\bm{\theta}),\text{diag}(\bm{\lambda}^{2})\bigr{)},$ (3.9)
where $\bm{\lambda}\in\mathbb{R}^{F\times T}$ is a vector that defines the scale of the output covariance. In practice, $\bm{\lambda}$ can either be fixed (e.g., $\bm{\lambda}=\lambda\mathbf{1}$ for some scalar $\lambda>0$), which is also the case in the present work, or learned as part of the model, depending on the modeling assumptions. Notice that during training, the decoder takes latent variables $\mathbf{z}$ sampled from the approximate posterior $q_{\bm{\phi}}(\mathbf{z}\mid\mathbf{x})$ and learns to approximate the likelihood $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$. During generation, it takes $\mathbf{z}$ sampled from the prior $p(\mathbf{z})$ and generates new samples ${\mathbf{x}}$. That is, the decoder bridges the latent space and the data space.
@@ 3.3 Conditional Variational Autoencoder (CVAE)
    $\mathcal{NN}_{enc}(\mathbf{x},\mathbf{c};\bm{\phi})=(\bm{\mu}_{\bm{\phi}}(\mathbf{x},\mathbf{c}),\text{diag}(\bm{\sigma}_{\bm{\phi}}^{2}(\mathbf{x},\mathbf{c}))).$ (3.22)
@@
    $p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z},\mathbf{c})=\mathcal{N}({\mathbf{x}}\mid\mathcal{NN}_{dec}(\mathbf{z},\mathbf{c};\bm{\theta})).$ (3.23)
@@ 3.4 Gaussian Mixture Model Conditional Variational Autoencoder (GMM-CVAE)
During training, the encoder network $\mathcal{NN}_{enc}(\mathbf{x},\mathbf{c};\bm{\phi})$ learns the conditioned mean and covariance $(\bm{\mu}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}),\bm{\Sigma}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}))$ of the variational posterior $q_{\bm{\phi}}(\mathbf{z}\mid\mathbf{x},\mathbf{c})$, which is defined as,

### v2

@@ Abstract
This work proposes variational autoencoders (VAEs) to predict a vehicle’s jerk signals from torque demand in the context of limited real-world drivetrain datasets. We implement both unconditional and conditional VAEs, trained on experimental data from two variants of a fully electric SUV with differing torque and drivetrain configurations. The VAEs synthesize jerk signals that capture characteristics from multiple drivetrain scenarios by leveraging the learned latent space. A performance comparison with baseline physics-based and hybrid models confirms the effectiveness of the VAEs, without requiring detailed system parametrization. Unconditional VAEs generate realistic jerk signals without prior system knowledge, while conditional VAEs enable the generation of signals tailored to specific torque inputs. This approach reduces the dependence on costly and time-intensive real-world experiments and extensive manual modeling. The results support the integration of generative models such as VAEs into drivetrain simulation pipelines, both for data augmentation and for efficient exploration of complex operational scenarios, with the potential to streamline validation and accelerate vehicle development.
@@ 1 Introduction
In automotive engineering, engineers must rely on datasets to accurately parameterize and validate the generalization capabilities of drivetrain simulation models. Ideally, these datasets should cover the variance of drivetrain hardware characteristics across the vehicle fleet and the broad variety of driving situations. However, the issue is that measurement collection in the real world can be very expensive [1]. Consequently, simulation engineers often have to work with sparse datasets that, for example, do not fully capture the variance in the jerk signal behavior across the vehicle fleet. A vehicle’s drivetrain encompasses the components that deliver power from the engine/motor to the wheels, while jerk refers to the rate of change of acceleration, a key factor in perceived ride comfort. Vehicles with different mileage or drivetrain components with slightly different clearances can exhibit significant differences in their jerk behavior. A sparse or poorly distributed dataset, therefore, limits the generalization ability of a drivetrain model.
@@
To mitigate this issue, synthetic measurements can be generated using non-generative methods that rely on models and their corresponding inputs to simulate new signals of interest [2]. Such a supervised approach to data augmentation can be greatly limited if the required input data for a physics-based or data-driven model are missing in the recorded measurement files or cannot be measured at all [3]. Another major drawback of this conventional approach is that it requires drivetrain engineers to deeply understand the underlying physical phenomena to determine which input data are necessary for a non-generative model to simulate additional valuable signals [4]. Furthermore, the more physically relevant input information is fed into the model for precise simulation results, the more accurate the simulation will likely be. This input-output imbalance necessitates a large number of input parameters (like torque and speed of the electric machine) to simulate a single particular output, such as vehicle acceleration.
@@
These limitations can be tackled with approaches from generative artificial intelligence (AI) [5, 6]. One such family of deep generative models includes unconditional architectures such as Variational Autoencoders (VAEs) [7, 8], which can be trained to generate new signals solely from a limited amount of available signals collected from physics-based, simulations, or a real system and does not require additional parametrization with respect to the system configuration. This self-supervised approach to data augmentation is purely data-driven. It eliminates the need for detailed domain expertise regarding the underlying physical principles of drivetrain simulation. While unconditional data augmentation can be highly advantageous for drivetrain engineers, it is also important to consider the value of conditional AI, for example, conditional VAEs (CVAEs) [9, 10, 11]. In this case, the generation of synthetic data is additionally conditioned on other inputs, similar to supervised data augmentation. This can significantly minimize the need for real-world on-road testing, making generative AI a powerful tool for addressing incomplete datasets.
@@
The outline of this paper is as follows. First, we introduce the drivetrain simulation problem in Section 2 and explain the different possible approaches to address it. Second, we present the theoretical background for VAEs in Section 3 and extensions that yield conditional VAEs. Section 4 presents the dataset and the challenges it imposes for VAE training. We then train unconditional and conditional VAEs to address the problem of jerk signal generation and measure their performance using different metrics. We also compare the performance of the trained VAEs with respect to a physics-based and a hybrid model. The manuscript concludes in Section 5.
@@ 2 Drivetrain simulation problem
A key performance indicator for quantifying vehicle drivability is the jerk, defined as the rate of change of acceleration. During abrupt positive changes in acceleration, known as tip-in load changes, the torsion in the driveshaft fluctuates, stimulating vibrations in the natural frequency range of the drivetrain (0-2 Hz and 8-12 Hz) [15]. These longitudinal oscillations negatively affect vehicle dynamics, leading to reduced drivability performance and driver discomfort [15, 16]. Although this issue is present in vehicles with combustion engines, it becomes more pronounced in electric vehicles due to their higher torque gradients [17]. The primary control variable for managing the jerk is torque demand, although it is also influenced by drivetrain parameters such as gear ratio, vehicle mass, and tire dynamics. Furthermore, parameters such as electric machine (ELM) speed and road conditions–referred to as driving scenario parameters–significantly impact jerk behavior.
@@
However, the accuracy of these models comes at the cost of complexity. Detailed physics-based models, which have been iteratively refined by experts over time, require deep domain knowledge and significant experience. Although effective, these traditional models are not always suitable for the fast-paced and disruptive development cycles of modern electric vehicles, where rapid iteration and efficiency are critical [18].
@@
Hybrid models combine the strengths of both data-driven and physics-based approaches [22, 36, 37, 38]. In the context of drivetrain simulation, they typically begin with a simplified physics-based model that captures the most critical physical phenomena. This ensures that the essential dynamics are grounded in physics. A data-driven correction term is added to account for any modeling discrepancies or to refine the model [39].
@@ 3.1 Transformation between time domain jerk signal and spectrogram
Projecting time-domain signals into the spectral domain (spectrograms) was chosen as it offers several advantages for processing with VAEs, especially those utilizing convolutional neural network (CNN) encoders and decoders. Spectrograms provide a rich time-frequency representation, allowing the model to learn and capture localized temporal and spectral patterns simultaneously. This can be more effective than using raw time-series data, as CNNs can more readily extract relevant features from the 2D structure of a spectrogram, potentially simplifying the network architecture and improving the ability to model complex signal characteristics.
@@
where $\operatorname{Re}(S(f,t))$ and $\operatorname{Im}(S(f,t))$ are the real and imaginary parts of $S(f,t)$, respectively.
To compress the dynamic range of the magnitude spectrogram, a logarithmic scaling is applied,
@@
This logarithmic scaling is a common practice in signal processing as it helps to enhance the details of lower-energy components in the spectrogram and makes the distribution of the input features more suitable for neural network training by preventing high-magnitude values from disproportionately influencing the learning process. For jerk signals, oscillations at lower frequencies can be particularly important for drivability perception, and log-scaling can aid in better representing these nuances.
@@ 3.2 Variational Autoencoder (VAE)
The hidden random generative process of $\mathbf{x}$ consists of two steps: First, the latent variable $\mathbf{z}$ is sampled from some prior distribution $p(\mathbf{z})$. This prior $p(\mathbf{z})$ is chosen to be a standard Gaussian distribution, $p(\mathbf{z})=\mathcal{N}(\mathbf{z}\mid\mathbf{0},\mathbf{I})$, where $\mathbf{I}$ is the identity matrix. This choice simplifies calculations, particularly for the regularization term in the objective function, and encourages a well-behaved latent space. Thus, $\mathbf{z}\sim p(\mathbf{z})$. Second, a value $\mathbf{x}$ is generated by some conditional distribution $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$, i.e. $\mathbf{x}\sim p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$. We aim to find $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ (the parameters $\bm{\theta}$), such that we can sample from $p(\mathbf{z})$ and with high probability, $\mathbf{x}\sim p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ follows the same distribution as the $\mathbf{x}$ in our dataset [9]. From a mathematical point of view, this can be achieved by maximizing the marginal log-likelihood of $\mathbf{x}$ (also termed the log evidence). The marginal log-likelihood is expressed as:
$\log p_{\bm{\theta}}(\mathbf{x})=\log\int p_{\bm{\theta}}(\mathbf{x},\mathbf{z})\,\text{d}\mathbf{z}=\log\int\,p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})p(\mathbf{z})\,\text{d}\mathbf{z}.$ (3.5)
Here, $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ represents the likelihood of observing data $\mathbf{x}$ given latent variables $\mathbf{z}$ and model parameters $\bm{\theta}$, while $p(\mathbf{z})$ is the prior over $\mathbf{z}$. Maximizing $\log p_{\bm{\theta}}(\mathbf{x})$ with respect to $\bm{\theta}$ inherently optimizes the parameters of the likelihood function $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$. This is because, to render the observed data $\mathbf{x}$ maximally probable under the integral, the likelihood $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ must assign high probabilities to $\mathbf{x}$ for latent variables $\mathbf{z}$ that are themselves likely under the prior $p(\mathbf{z})$. Thus, this process compels the model to learn a likelihood function that effectively explains the generation of the observed data. This marginal likelihood $p_{\bm{\theta}}(\mathbf{x})$ is challenging to compute because:
  • The latent variable $\mathbf{z}$ may be high-dimensional, making Monte Carlo-based integration over $\mathbf{z}$ practically infeasible.
  • While the likelihood $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ (parameterized by the decoder) is typically tractable for a given $\mathbf{z}$, the integration $\int p(\mathbf{z})\,p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})\,\text{d}\mathbf{z}$ for computing the marginal likelihood $p_{\bm{\theta}}(\mathbf{x})$ is often intractable, especially when $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$ is a complex function (e.g., a deep neural network) or when no closed-form solution for the integral exists.
@@
$(\bm{\mu}_{\bm{\phi}}(\mathbf{x}),\bm{\sigma}_{\bm{\phi}}(\mathbf{x}))={\text{NN}}_{\text{enc}}(\mathbf{x};\bm{\phi}),$ (3.8)
where $\text{NN}_{\text{enc}}$ is the encoder neural network parametrized in $\bm{\phi}$, that branches into the two outputs $(\bm{\mu}_{\bm{\phi}}(\mathbf{x}),\bm{\sigma}_{\bm{\phi}}(\mathbf{x}))$. The parameter set $\bm{\phi}$ includes all weights and biases necessary to produce both $\bm{\mu}_{\bm{\phi}}(\mathbf{x})$ and $\bm{\sigma}_{\bm{\phi}}(\mathbf{x})$. The encoder bridges the data space and the latent space.
Decoder: It approximates the likelihood $p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z})$ with a neural network $\text{NN}_{\mathrm{dec}}(\mathbf{z};\bm{\theta})$, parameterized by $\bm{\theta}$, to output the mean of a Gaussian distribution
$p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z})=\mathcal{N}\bigl{(}{\mathbf{x}}\mid\text{NN}_{\mathrm{dec}}(\mathbf{z};\bm{\theta}),\text{diag}(\bm{\lambda}^{2})\bigr{)},$ (3.9)
where $\bm{\lambda}\in\mathbb{R}^{F\times T}$ is a vector that defines the scale of the output covariance. In practice, $\bm{\lambda}$ can either be fixed (e.g., $\bm{\lambda}=\lambda\mathbf{1}$ for some scalar $\lambda>0$), or learned as part of the model, depending on the modeling assumptions. In this study, we set $\lambda$ to a constant value of $1.0$ for all output dimensions. This choice makes the negative log-likelihood reconstruction term in the ELBO (Equation 3.17) equivalent to a scaled Mean Squared Error plus a constant, a common practice that encourages the decoder to learn the mean of the data distribution effectively. While more complex output distributions are possible, this choice provides a balance between model expressiveness and computational tractability for many types of continuous data. Notice that during training, the decoder takes latent variables $\mathbf{z}$ sampled from the approximate posterior $q_{\bm{\phi}}(\mathbf{z}\mid\mathbf{x})$ and learns to approximate the likelihood $p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z})$. During generation, it takes $\mathbf{z}$ sampled from the prior $p(\mathbf{z})$ and generates new samples ${\mathbf{x}}$. That is, the decoder bridges the latent space and the data space.
@@ 3.3 Conditional Variational Autoencoder (CVAE)
    ${\text{NN}}_{\text{enc}}(\mathbf{x},\mathbf{c};\bm{\phi})=(\bm{\mu}_{\bm{\phi}}(\mathbf{x},\mathbf{c}),\text{diag}(\bm{\sigma}_{\bm{\phi}}^{2}(\mathbf{x},\mathbf{c}))).$ (3.22)
@@
    $p_{\bm{\theta}}({\mathbf{x}}\mid\mathbf{z},\mathbf{c})=\mathcal{N}({\mathbf{x}}\mid{\text{NN}}_{\text{dec}}(\mathbf{z},\mathbf{c};\bm{\theta}),\text{diag}(\bm{\lambda}^{2})).$ (3.23)
@@ 3.4 Gaussian Mixture Model Conditional Variational Autoencoder (GMM-CVAE)
During training, the encoder network ${\text{NN}}_{enc}(\mathbf{x},\mathbf{c};\bm{\phi})$ learns the conditioned mean and covariance $(\bm{\mu}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}),\bm{\Sigma}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}))$ of the variational posterior $q_{\bm{\phi}}(\mathbf{z}\mid\mathbf{x},\mathbf{c})$, which is defined as,
@@
Here, the notation $\mathcal{N}(\mathbf{z}\mid\bm{\mu},\bm{\Sigma})$ refers to the probability density function of the Gaussian distribution evaluated at a sample $\mathbf{z}$. The responsibilities $\gamma_{k}(\mathbf{x},\mathbf{c})$ effectively denote the posterior probability of the latent variable $\mathbf{z}$ (associated with input $\mathbf{x}$ under condition $\mathbf{c}$) belonging to the $k$-th Gaussian component of the mixture. During training, to sample from $q_{\bm{\phi}}(\mathbf{z}\mid\mathbf{x},\mathbf{c})$, a component $k$ is typically chosen according to these responsibilities (interpreted as $P(k\mid\mathbf{x},\mathbf{c})$), and then $\mathbf{z}$ is sampled from $\mathcal{N}(\mathbf{z}\mid\bm{\mu}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}),\bm{\Sigma}_{\bm{\phi},k}(\mathbf{x},\mathbf{c}))$.